# Bayesian Optimization with Gaussian Processes to Accelerate Stationary Point Searches

### 11 Mar 2026 v1

@@ 9.1  Muller-Brown Potential
The Muller-Brown surface [70] (Figure 11) is a standard benchmark for saddle point searches, with three minima and two saddle points connected by curved MEPs. Figure LABEL:fig:rpc_dimer_convergence compares the standard dimer, GP-dimer, and OTGPD on a representative system: the OTGPD converges with the fewest oracle calls, while the GP-dimer shows oscillations that the OT-GP extensions suppress. The GP surface after convergence closely matches the true PES within the trust region (Figure 10), but diverges outside it, which is the expected behavior of a local surrogate. The NEB path connecting minima A and B through saddle S2 is shown in Figure 11.
@@ 9.3  PET-MAD Molecular System
The PET-MAD universal potential [64], trained on the MAD dataset [65], provides a realistic test beyond toy surfaces. Figure 19 shows the convergence of GP-accelerated minimization on a molecular system, where the maximum per-atom force drops below the convergence threshold with significantly fewer oracle evaluations than the unaccelerated optimizer.
@@
For NEB, a 9-atom cycloaddition system ($\text{C}_{2}\text{H}_{4}+\text{N}_{2}\text{O}$, 27 degrees of freedom, 36 inverse-distance features) on the PET-MAD surface illustrates how the GP-NEB variants scale to molecular reactions. Figure 20 compares classical NEB, AIE, and OIE on this system. A relaxed convergence threshold of 0.5 eV/Å is used here for demonstration purposes; tighter tolerances are straightforward but require more evaluations and are better explored in a dedicated benchmarking study. Classical NEB converges in 132 oracle evaluations. The AIE variant converges in 72 evaluations, showing that the surrogate accelerates convergence even when all images are evaluated each cycle. The OIE variant evaluates only one image per cycle, selected by maximum predictive variance, and converges in 36 evaluations, a $\sim 3.7\times$ reduction over classical NEB. Figure 21 shows the energy profiles along the converged MEP; all three variants recover the same barrier ($\sim 1.2$ eV) and exothermic product basin ($\Delta E\approx-0.43$ eV), confirming that the surrogate does not distort the path. Figure 22 shows the reaction valley projection [33] comparing the standard NEB and AIE saddle points; both paths lie within the low-variance region ($\sigma^{2}=0.046$) where the surrogate is well-trained, and the molecular snapshots along the bottom illustrate the structural progression from reactant to product.
Figure 20: Convergence of GP-NEB variants on a 9-atom cycloaddition (PET-MAD surface, 27 DOF). Climbing image force versus oracle evaluations on a logarithmic scale. Classical NEB (132 calls), AIE (72 calls), and OIE (36 calls). Dashed line shows the convergence threshold.
Figure 21: Energy profiles along the converged MEP for the cycloaddition system on the PET-MAD surface. All three NEB variants recover the same barrier ($\sim 1.2$ eV) and exothermic product basin ($\Delta E\approx-0.43$ eV). The AIE profile is slightly shifted near the saddle region but converges to the same endpoints.
Figure 22: Reaction valley projection [33] of the NEB paths on the PET-MAD surface. Reaction progress $s$ and orthogonal deviation $d$ are projected RMSD coordinates. Standard NEB (blue stars) and GP-NEB AIE (orange stars) saddle points are compared; both lie near the true saddle (black square). Dashed contours show GP variance ($\sigma^{2}$); the paths remain within the low-variance ($\sigma^{2}=0.046$) region where the surrogate is reliable. Bottom: molecular snapshots at key points along the reaction coordinate.
@@ 11 Supporting Information ignoreheading:appendix
The following supplementary material provides additional details on marginal likelihood landscapes, hyperparameter defaults, and code implementation details.
Appendix A Hyperparameter Defaults
@@
Appendix B Connection to Code: chemgp-core
@@
B.1  Kernel Evaluation and Analytical Derivative Blocks
@@
B.2  Covariance Matrix Assembly and GP Training
@@ B.3  Code Listings
Appendix C Mathematical Derivations
C.1  Rigid-Body Mode Basis Construction
The projection of rigid-body modes (Eq. LABEL:eq:rotation_removal in the main text) requires an orthonormal basis $\{\mathbf{u}_{k}\}_{k=1}^{6}$ spanning the 6 external degrees of freedom. For a molecule with $N$ atoms and Cartesian coordinates $\mathbf{x}=(x_{1},y_{1},z_{1},\ldots,x_{N},y_{N},z_{N})^{T}\in\mathbb{R}^{3N}$, the six basis vectors correspond to three translations and three infinitesimal rotations.
@@ C.1  Rigid-Body Mode Basis Construction
$\mathbf{t}_{x}$ $=(1,0,0,1,0,0,\ldots,1,0,0)^{T},$ (C.1)
$\mathbf{t}_{y}$ $=(0,1,0,0,1,0,\ldots,0,1,0)^{T},$ (C.2)
$\mathbf{t}_{z}$ $=(0,0,1,0,0,1,\ldots,0,0,1)^{T}.$ (C.3)
@@
$\mathbf{r}_{x}$ $=(0,z_{1}-z_{0},-(y_{1}-y_{0}),\ldots,0,z_{N}-z_{0},-(y_{N}-y_{0}))^{T},$ (C.4)
$\mathbf{r}_{y}$ $=(-(z_{1}-z_{0}),0,x_{1}-x_{0},\ldots,-(z_{N}-z_{0}),0,x_{N}-x_{0})^{T},$ (C.5)
$\mathbf{r}_{z}$ $=(y_{1}-y_{0},-(x_{1}-x_{0}),0,\ldots,y_{N}-y_{0},-(x_{N}-x_{0}),0)^{T}.$ (C.6)
@@
$\mathbf{u}_{1}$ $=\frac{\mathbf{t}_{x}}{\|\mathbf{t}_{x}\|},$ (C.7)
$\mathbf{u}_{2}$ $=\frac{\mathbf{t}_{y}-(\mathbf{t}_{y}\cdot\mathbf{u}_{1})\mathbf{u}_{1}}{\|\mathbf{t}_{y}-(\mathbf{t}_{y}\cdot\mathbf{u}_{1})\mathbf{u}_{1}\|},$ (C.8)
$\mathbf{u}_{3}$ $=\frac{\mathbf{t}_{z}-\sum_{j=1}^{2}(\mathbf{t}_{z}\cdot\mathbf{u}_{j})\mathbf{u}_{j}}{\|\mathbf{t}_{z}-\sum_{j=1}^{2}(\mathbf{t}_{z}\cdot\mathbf{u}_{j})\mathbf{u}_{j}\|},$ (C.9)
$\mathbf{u}_{k+3}$ $=\frac{\mathbf{r}_{k}-\sum_{j=1}^{k+2}(\mathbf{r}_{k}\cdot\mathbf{u}_{j})\mathbf{u}_{j}}{\|\mathbf{r}_{k}-\sum_{j=1}^{k+2}(\mathbf{r}_{k}\cdot\mathbf{u}_{j})\mathbf{u}_{j}\|},\quad k\in\{x,y,z\}.$ (C.10)
In practice, the translation vectors are already orthogonal (and of equal norm $\sqrt{N}$), so the Gram-Schmidt process only needs to orthogonalize the rotation vectors against the translations and against each other. This basis is computed once per molecular geometry and cached; the projection (Eq. LABEL:eq:rotation_removal) then costs $\mathcal{O}(N)$ per step.
C.2  Hyperparameter Oscillation Detection
@@
$O_{j}(t)=\begin{cases}1,&\text{if }(\theta_{j}(t)-\theta_{j}(t-1))(\theta_{j}(t-1)-\theta_{j}(t-2))<0,\\ 0,&\text{otherwise}.\end{cases}$ (C.11)
@@
C.3  Kernel Block Structure
@@
$k_{ee}$ $=k(\phi(\mathbf{x}_{1}),\phi(\mathbf{x}_{2}))$ (C.13)
$\ k_{ef}$ $=\frac{\partial k}{\partial\mathbf{x}_{2}}=\left(\frac{\partial k}{\partial\boldsymbol{\phi}}\right)^{T}\frac{\partial\boldsymbol{\phi}}{\partial\mathbf{x}_{2}}=\mathbf{J}_{2}^{T}\frac{\partial k}{\partial\boldsymbol{\phi}}$ (C.14)
$\ k_{fe}$ $=\frac{\partial k}{\partial\mathbf{x}_{1}}=\mathbf{J}_{1}^{T}\frac{\partial k}{\partial\boldsymbol{\phi}}$ (C.15)
$\ k_{ff}$ $=\frac{\partial^{2}k}{\partial\mathbf{x}_{1}\partial\mathbf{x}_{2}^{T}}=\mathbf{J}_{1}^{T}\left(\frac{\partial^{2}k}{\partial\boldsymbol{\phi}\partial\boldsymbol{\phi}^{T}}-2\frac{\partial k}{\partial\boldsymbol{\phi}}\otimes\frac{\partial\boldsymbol{\phi}}{\partial\mathbf{x}_{2}}\right)\mathbf{J}_{2}$ (C.16)
@@
C.4  Earth Mover’s Distance for Trust Regions
@@
$d_{\text{EMD}}^{(t)}=\min_{\pi\in\Pi_{t}}\sum_{i\in\mathcal{I}_{t}^{(1)}}\sum_{j\in\mathcal{I}_{t}^{(2)}}c_{ij}^{(t)}\pi_{ij}$ (C.17)
@@
$d_{\text{EMD}}(\mathbf{x}_{1},\mathbf{x}_{2})=\sum_{t}w_{t}\cdot d_{\text{EMD}}^{(t)}$ (C.18)
@@
C.5  OIE Acquisition Criterion Derivation
@@
$\alpha(\mathbf{R}_{i})=|\mathbf{F}_{i}^{\text{NEB}}|+\kappa\cdot\sigma_{\perp}(\mathbf{R}_{i},\boldsymbol{\tau}_{i})$ (C.19)
@@
$|F_{d}^{\text{true}}-F_{d}^{\text{GP}}|\leq\kappa\cdot\sigma(F_{d}^{\text{GP}}),\quad\kappa=\sqrt{2\log(1/\delta)}$ (C.20)

### 16 Mar 2026 v2

@@ 9  Add $(\mathbf{x},V,\nabla V)$ to $\mathcal{D}$;
Figure 9: Convergence comparison of the standard dimer, GP-dimer, and OT-GP dimer (OTGPD) on a molecular system (C3H5 allyl radical, 8 atoms, 24 DOF) via eOn serve mode. The maximum per-atom force (eV/Å) is plotted against oracle evaluations on a logarithmic scale. The OTGPD variant reaches the convergence threshold (gray dashed, 0.1 eV/Å) with the fewest oracle calls; the GP-dimer shows oscillations from surrogate retraining instabilities that the OT-GP extensions suppress.
@@ 9.1  Muller-Brown Potential
The Muller-Brown surface [70] (Figure 12) is a standard benchmark for saddle point searches, with three minima and two saddle points connected by curved MEPs. Because the surface is two-dimensional, the GP behavior can be visualized directly: Figure 3 shows how the surrogate evolves from a crude approximation with few training points to a faithful replica of the true PES as data accumulates. The variance landscape (Figure 4) guides the active learning criterion, concentrating evaluations near the reaction path. The GP surface closely matches the true PES within the trust region (Figure 11), but diverges outside it, which is the expected behavior of a local surrogate. The NEB path connecting minima A and B through saddle S2 is shown in Figure 12.
@@ 9.3  PET-MAD Molecular System
The PET-MAD universal potential [64], trained on the MAD dataset [65], provides a realistic test beyond toy surfaces where the inverse-distance kernel operates on physical interatomic distances. For saddle point searches, Figure 9 compares the standard dimer, GP-dimer, and OTGPD on the C3H5 allyl radical (8 atoms, 24 DOF): the OTGPD converges with the fewest oracle calls, while the GP-dimer shows oscillations that the OT-GP extensions suppress. Figure 20 shows the convergence of GP-accelerated minimization on a molecular system, where the maximum per-atom force drops below the convergence threshold with significantly fewer oracle evaluations than the unaccelerated optimizer.
@@
For NEB, a 9-atom cycloaddition system ($\text{C}_{2}\text{H}_{4}+\text{N}_{2}\text{O}$, 27 degrees of freedom, 36 inverse-distance features) on the PET-MAD surface illustrates how the GP-NEB variants scale to molecular reactions. Figure 21 compares classical NEB, AIE, and OIE on this system. A relaxed convergence threshold of 0.5 eV/Å is used here for demonstration purposes; tighter tolerances are straightforward but require more evaluations and are better explored in a dedicated benchmarking study. Classical NEB converges in 102 oracle evaluations. The AIE variant converges in 42 evaluations, showing that the surrogate accelerates convergence even when all images are evaluated each cycle. The OIE variant evaluates only one image per cycle, selected by maximum predictive variance, and converges in 44 evaluations, a $\sim 2.3\times$ reduction over classical NEB. Figure 22 shows the energy profiles along the converged MEP; all three variants recover the same barrier and exothermic product basin, confirming that the surrogate does not distort the path. Figure 23 shows the reaction valley projection [33] comparing the standard NEB and AIE saddle points; both paths lie within the low-variance region where the surrogate is well-trained, and the molecular snapshots along the bottom illustrate the structural progression from reactant to product.
Figure 21: Convergence of GP-NEB variants on a 9-atom cycloaddition (PET-MAD surface, 27 DOF). Climbing image force versus oracle evaluations on a logarithmic scale. Classical NEB (102 calls), AIE (42 calls), and OIE (44 calls). Dashed line shows the convergence threshold.
Figure 22: Energy profiles along the converged MEP for the cycloaddition system on the PET-MAD surface. All three NEB variants recover the same barrier and exothermic product basin. The AIE profile is slightly shifted near the saddle region but converges to the same endpoints.
Figure 23: Reaction valley projection [33] of the NEB paths on the PET-MAD surface. Reaction progress $s$ and orthogonal deviation $d$ are projected RMSD coordinates. Standard NEB (blue stars) and GP-NEB AIE (orange stars) saddle points are compared; both lie near the true saddle (black square). Dashed contours show GP variance ($\sigma^{2}$); the paths remain within the low-variance region where the surrogate is reliable. Bottom: molecular snapshots at key points along the reaction coordinate.
@@ 11 Supporting Information ignoreheading:appendix
The following supplementary material provides additional details on algorithm flowcharts, marginal likelihood landscapes, hyperparameter defaults, and code implementation details.
.1  Algorithm Flowcharts
The following flowcharts visualize the classical and GP-accelerated versions of the dimer method and NEB. Each pair shows the unmodified algorithm (left) alongside its GP-accelerated counterpart (right).
Figure 24: Classical dimer (left) and GP-dimer (right) flowcharts. The classical version evaluates the true PES at every rotation step; the GP-accelerated version performs rotations and translations on the surrogate surface, querying the oracle only when trust radius violations occur or GP-level convergence is achieved.
Figure 25: Classical NEB (left) and GP-NEB (right) flowcharts. The classical version optimizes all images via L-BFGS; the GP-NEB variant uses OIE acquisition to select a single image per outer loop iteration, minimizing oracle queries.
Appendix A Marginal Likelihood Landscape
Figure 26: MAP-regularized negative log marginal likelihood landscape on the LEPS surface. The filled contours show the NLL as a function of the log-hyperparameters $\ln\sigma^{2}$ and $\ln\theta$, computed on a 40 $\times$ 40 grid from 5 training points near the reactant. The dashed black contours show the gradient norm. The coral star marks the MAP optimum that SCG converges to. Regions where the Cholesky factorization fails (hyperparameters that produce a non-positive-definite covariance matrix) are masked.
Appendix B Hyperparameter Defaults
@@
Appendix C Connection to Code: chemgp-core
@@
C.1  Kernel Evaluation and Analytical Derivative Blocks
@@
C.2  Covariance Matrix Assembly and GP Training
@@ C.3  Code Listings
Appendix D Mathematical Derivations
D.1  Rigid-Body Mode Basis Construction
The projection of rigid-body modes requires an orthonormal basis $\{\mathbf{u}_{k}\}_{k=1}^{6}$ spanning the 6 external degrees of freedom. For a molecule with $N$ atoms and Cartesian coordinates $\mathbf{x}=(x_{1},y_{1},z_{1},\ldots,x_{N},y_{N},z_{N})^{T}\in\mathbb{R}^{3N}$, the six basis vectors correspond to three translations and three infinitesimal rotations.
@@ D.1  Rigid-Body Mode Basis Construction
$\mathbf{t}_{x}$ $=(1,0,0,1,0,0,\ldots,1,0,0)^{T},$ (D.1)
$\mathbf{t}_{y}$ $=(0,1,0,0,1,0,\ldots,0,1,0)^{T},$ (D.2)
$\mathbf{t}_{z}$ $=(0,0,1,0,0,1,\ldots,0,0,1)^{T}.$ (D.3)
@@
$\mathbf{r}_{x}$ $=(0,z_{1}-z_{0},-(y_{1}-y_{0}),\ldots,0,z_{N}-z_{0},-(y_{N}-y_{0}))^{T},$ (D.4)
$\mathbf{r}_{y}$ $=(-(z_{1}-z_{0}),0,x_{1}-x_{0},\ldots,-(z_{N}-z_{0}),0,x_{N}-x_{0})^{T},$ (D.5)
$\mathbf{r}_{z}$ $=(y_{1}-y_{0},-(x_{1}-x_{0}),0,\ldots,y_{N}-y_{0},-(x_{N}-x_{0}),0)^{T}.$ (D.6)
@@
$\mathbf{u}_{1}$ $=\frac{\mathbf{t}_{x}}{\|\mathbf{t}_{x}\|},$ (D.7)
$\mathbf{u}_{2}$ $=\frac{\mathbf{t}_{y}-(\mathbf{t}_{y}\cdot\mathbf{u}_{1})\mathbf{u}_{1}}{\|\mathbf{t}_{y}-(\mathbf{t}_{y}\cdot\mathbf{u}_{1})\mathbf{u}_{1}\|},$ (D.8)
$\mathbf{u}_{3}$ $=\frac{\mathbf{t}_{z}-\sum_{j=1}^{2}(\mathbf{t}_{z}\cdot\mathbf{u}_{j})\mathbf{u}_{j}}{\|\mathbf{t}_{z}-\sum_{j=1}^{2}(\mathbf{t}_{z}\cdot\mathbf{u}_{j})\mathbf{u}_{j}\|},$ (D.9)
$\mathbf{u}_{k+3}$ $=\frac{\mathbf{r}_{k}-\sum_{j=1}^{k+2}(\mathbf{r}_{k}\cdot\mathbf{u}_{j})\mathbf{u}_{j}}{\|\mathbf{r}_{k}-\sum_{j=1}^{k+2}(\mathbf{r}_{k}\cdot\mathbf{u}_{j})\mathbf{u}_{j}\|},\quad k\in\{x,y,z\}.$ (D.10)
In practice, the translation vectors are already orthogonal (and of equal norm $\sqrt{N}$), so the Gram-Schmidt process only needs to orthogonalize the rotation vectors against the translations and against each other. This basis is computed once per molecular geometry and cached; the projection then costs $\mathcal{O}(N)$ per step.
D.2  Hyperparameter Oscillation Detection
@@
$O_{j}(t)=\begin{cases}1,&\text{if }(\theta_{j}(t)-\theta_{j}(t-1))(\theta_{j}(t-1)-\theta_{j}(t-2))<0,\\ 0,&\text{otherwise}.\end{cases}$ (D.11)
@@
D.3  Kernel Block Structure
@@
$k_{ee}$ $=k(\phi(\mathbf{x}_{1}),\phi(\mathbf{x}_{2}))$ (D.13)
$\ k_{ef}$ $=\frac{\partial k}{\partial\mathbf{x}_{2}}=\left(\frac{\partial k}{\partial\boldsymbol{\phi}}\right)^{T}\frac{\partial\boldsymbol{\phi}}{\partial\mathbf{x}_{2}}=\mathbf{J}_{2}^{T}\frac{\partial k}{\partial\boldsymbol{\phi}}$ (D.14)
$\ k_{fe}$ $=\frac{\partial k}{\partial\mathbf{x}_{1}}=\mathbf{J}_{1}^{T}\frac{\partial k}{\partial\boldsymbol{\phi}}$ (D.15)
$\ k_{ff}$ $=\frac{\partial^{2}k}{\partial\mathbf{x}_{1}\partial\mathbf{x}_{2}^{T}}=\mathbf{J}_{1}^{T}\left(\frac{\partial^{2}k}{\partial\boldsymbol{\phi}\partial\boldsymbol{\phi}^{T}}-2\frac{\partial k}{\partial\boldsymbol{\phi}}\otimes\frac{\partial\boldsymbol{\phi}}{\partial\mathbf{x}_{2}}\right)\mathbf{J}_{2}$ (D.16)
@@
D.4  Earth Mover’s Distance for Trust Regions
@@
$d_{\text{EMD}}^{(t)}=\min_{\pi\in\Pi_{t}}\sum_{i\in\mathcal{I}_{t}^{(1)}}\sum_{j\in\mathcal{I}_{t}^{(2)}}c_{ij}^{(t)}\pi_{ij}$ (D.17)
@@
$d_{\text{EMD}}(\mathbf{x}_{1},\mathbf{x}_{2})=\sum_{t}w_{t}\cdot d_{\text{EMD}}^{(t)}$ (D.18)
@@
D.5  OIE Acquisition Criterion Derivation
@@
$\alpha(\mathbf{R}_{i})=|\mathbf{F}_{i}^{\text{NEB}}|+\kappa\cdot\sigma_{\perp}(\mathbf{R}_{i},\boldsymbol{\tau}_{i})$ (D.19)
@@
$|F_{d}^{\text{true}}-F_{d}^{\text{GP}}|\leq\kappa\cdot\sigma(F_{d}^{\text{GP}}),\quad\kappa=\sqrt{2\log(1/\delta)}$ (D.20)